OpenStudy (imtant):

Is it right? (2.7+0.55/1+0.55)×1.0Mg/m3×9.81m/sec2=20.56kN/m3

OpenStudy (anonymous):

is what right?

OpenStudy (imtant):

\[\left( 2.7+0.55 /1+0.55\right) \times 1.0Mg/m^{3} \times 9.81m/\sec^{2}=20.56 kN/m^{3}\]

OpenStudy (agent0smith):

I have no idea what you're trying to do.

OpenStudy (imtant):

I meant that,is it the correct answer to get in kN/m^3?